A harmonic wave is traveling along a rope. The oscillator that generates the wave completes 40.0 vibrations in 36.0 s. A given maximum of the wave travels 395 cm along the rope in a time period of 10.0 s. What is the wavelength?
1
Expert's answer
2014-01-15T10:22:37-0500
The wavelength λ of a sinusoidal waveform travelling at constant speed v is given by:
λ = v / f,
where f is the wave's frequency.
Using the data given:
f = 40.0 / 36.0 = 1.11 Hz
As 395 cm = 3.95 m:
v = 3.95 / 10.0 = 0.395 m/s
So, the wavelength λ = 0.395 / 1.11 = 0.356 m.
Answer: 0.356 m.